# Orthogonal Howe duality and dynamical (split) symmetric pairs

### v2

@@ Appendix A Technical lemmas
Denote by $D_{i}$ the operator $[D_{i}f](\mu):=\xi_{i}(\mu)f_{i}(\mu+2)$. For any rational functions $f_{1}(\mu),\ldots,f_{n}(\mu)$, consider the matrix
@@
We prove the statement by induction. The case $n=1$ is obvious. For the induction step, assume that $f_{1}(\lambda)\neq 0$; if there is $j$ such that $\xi_{j}(\mu)=\xi_{1}(\mu)$ then assume $j=2$. Let us divide the $i$-th row of $A$ by $[D_{1}^{i-1}f_{1}](\lambda)$ for every $i$. The matrix becomes
@@
where $[\tilde{D}_{i}f](\lambda):=\frac{\xi_{i}(\lambda)}{\xi_{1}(\lambda)}f(\lambda+2)$ and $\tilde{f}_{i}(\lambda):=\frac{f_{i}(\lambda)}{f_{1}(\lambda)}$, and its determinant is still zero. Additionally, let us subtract $(i-1)$-th row from the $i$-th one for every $i$. Then we get the matrix
@@
for every $k$ and sufficiently positive $\mu$. In particular, the system given by the matrix (A.1) with the same meaning of notations has a nonzero solution, hence its determinant zero for infinitely many $\mu$. Since the determinant is a rational function in $\mu$, it is zero for all $\mu$. Therefore, by A.1, we get that either one of the functions $f_{i}(\mu)$ is zero or there are $i,j$ such that $\xi_{i}(\mu)=\xi_{j}(\mu)$ (equivalently, $x_{i}=-x_{j}$) and $f_{i}(\mu)=c_{i}f_{j}(\mu)$. In the former case we can apply the induction assumption. As for the latter, denote by
@@
for all sufficiently positive integral $\mu$; now there are $n-1$ rational functions. Using property (7.8), we get for positive integral $\mu$
@@ Appendix B Calculations in rank two
In this appendix, we show the braid relation Theorem 7.18.
Let $\mathfrak{g}$ be a simple Lie algebra of rank $2$, i.e. $\mathfrak{sl}_{3}$, $\mathfrak{sp}_{4}$, or $\mathfrak{g}_{2}$. Let $\alpha$ (resp. $\beta$) be the short root (resp. the long root), and $\omega_{1}$ (resp. $\omega_{2}$) be the dual fundamental weight. Introduce coordinates on $\mathfrak{h}^{*}$ by $\lambda=\lambda_{1}\omega_{1}+\lambda_{2}\omega_{2}$.
@@
is satisfied if and only if the operators $\{A_{\alpha}^{\mathfrak{k}}(\lambda)\}$ satisfy the shifted one of Theorem 7.18.
@@
To prove Proposition 3.13 and Theorem 7.18, we perform explicit calculations to verify that the operators $R_{\alpha}$ and $R_{\beta}$ satisfy the braid relations in specific integrable $\mathfrak{k}$-modules $\{X_{i}\}$, such that any other irreducible integrable module is a direct $\mathfrak{k}$-summand of $X_{i}\otimes V$ for suitable $i$ and a finite-dimensional $\mathfrak{g}$-representation $V$. These calculations, along with arguments in the proof of Proposition 3.13 and Theorem 7.18, imply the braid relations in general integrable $\mathfrak{k}$-modules.
We will also prove 7.24 case-by-case. Denote by
@@ B.1.1. Braid relation in special cases
Recall the tensor property $\Delta\tilde{s}^{\mathfrak{k}}_{\alpha}=\tilde{s}^{\mathfrak{k}}_{\alpha}\otimes\tilde{s}^{\mathfrak{g}}_{\alpha}$. By the same arguments as in the proof of Proposition 3.13 or Theorem 7.18, it is enough to show this relation for $X=V(1)$ and $V$ an arbitrary $\mathfrak{sl}_{3}$-representation $V$. For the latter, it is enough to show it for the vector representation $\mathbb{C}^{3}$, as any $\mathfrak{sl}_{3}$-representation is a subrepresentation of a suitable tensor product of $\mathbb{C}^{3}$. Both cases can be verified directly.
@@ B.2. Type $C_{2}$
We can choose a basis of $\mathfrak{g}$ satisfying the assumptions of Section 4 as follows:
@@ B.2.1. Braid relation in special cases
The proof of Proposition 3.13 is similar to the dynamical case.
B.2.2. Proof of 7.24
@@
Therefore, the dynamical braid relation (B.9) for $S$-operators follows from the non-dynamical one of Proposition 3.13.
@@ B.3. Type $G_{2}$
One can check that $\{J_{i}\}$ and $\{J^{\prime}_{i}\}$ satisfy the $\mathfrak{so}_{3}$-relations (B.5) and $[J_{i},J^{\prime}_{k}]=0$ for every $i,k$. In particular, we obtain that $\mathfrak{k}:=\mathfrak{g}_{2}^{\theta}=\mathfrak{so}_{3}\oplus\mathfrak{so}_{3}$. In these generators, we have
@@ B.3.1. Braid relation in special cases
Then we proceed as in Section B.1 for both dynamical braid relation of Theorem 7.18 and non-dynamical one Proposition 3.13.
B.3.2. Proof of 7.24
@@
Denote by $b_{\alpha}^{\mathfrak{so}_{3}},b_{\beta}^{\mathfrak{so}_{3}}$ the type $A_{2}$ generators of Section B.1 and by $\Delta$ the coproduct on $\mathfrak{so}_{3}$. Then we have
@@ C.1. Motivation from dynamical $R$-matrix
ˇ$\check{\mathcal{R}}_{V,W}(\lambda):=J_{WV}(\lambda)^{-1}\circ P\circ J_{VW}(\lambda),$
@@
(ˇ$\left(\check{\mathcal{R}}_{V,W}(\lambda)\otimes\mathrm{id}_{U}\right)\circ\left(\mathrm{id}_{V}\otimes\check{\mathcal{R}}_{U,W}(\lambda-h|_{V})\right)\circ\left(\check{\mathcal{R}}_{U,V}(\lambda)\otimes\mathrm{id}_{W}\right).$
@@ C.2. Background on quantum $K$-matrix
If we write ˇ$\check{\mathcal{R}}_{q}:=P\circ\mathcal{R}_{q}$, then the $2$-tensor reflection equation becomes
@@
Recall that $-w_{0}$ permutes the simple roots, inducing an automorphism of the Dynkin diagram. Let $\sigma$ be the order two automorphism [^10] In general, $\sigma=\theta\theta_{0}$, where $\theta$ is the diagram automorphism in the data of the Satake diagram and $\theta_{0}$ is the automorphism of the Dynkin diagram of $\mathfrak{g}$ induced by $-w_{0}$, in this paper we concentrate on the split case – so $\theta=\mathrm{id}$ and $\sigma=\theta_{0}$. of $U_{q}(\mathfrak{g})$ such that $\sigma(E_{\alpha})=E_{-w_{0}(\alpha)}$ and $\sigma(F_{\alpha})=F_{-w_{0}(\alpha)}$ for $\alpha\in R$. The $1$-tensor universal $K$-matrix satisfies the following [Kolb-Kmatrixbraided, equation 1.1]
@@ C.3. Definition of $q=1$ $K$-matrix
Let $w_{0}$ denote the longest element in $W$. The operator $-w_{0}$ permutes simple roots, inducing an automorphism of the Dynkin diagram [^11] This automorphism may be $\mathrm{id}$, which is the case if $-1\in W\subset GL(\mathfrak{h}^{*})$. which we denote by $\sigma$. The Serre-presentation of $\mathfrak{g}$ implies that $\sigma$ determines an associative algebra automorphism of $\mathrm{U}(\mathfrak{g})$, denoted $\mathrm{Ad}_{\sigma}$ such that $\mathrm{Ad}_{\sigma}(x_{\alpha_{i}})=x_{\alpha_{\sigma(i)}}$, for $x\in\{f,h,e\}$ and $\alpha_{i}\in R$.
@@
As in Remark 3.2, let $G$ be the simply connected group with the Lie algebra $\mathfrak{g}$. For $g\in G$, we write $\mathrm{Ad}_{g}$ for the automorphism of $\mathfrak{g}$ defined by $x\mapsto gxg^{-1}$. The operators $\widetilde{w}$ can be viewed as elements in $G$ and thus act by conjugation on $\mathfrak{g}$. In particular, one has
@@
When $q=1$, we have $\mathcal{R}_{q=1}=1\otimes 1$ and ˇ$\check{\mathcal{R}}_{q=1}=P$. Therefore both sides of the $2$-tensor reflection equation become $\mathcal{K}_{q=1}\otimes\mathcal{K}_{q=1}$. Since $\Delta(\widetilde{w_{0}})=\widetilde{w_{0}}\otimes\widetilde{w_{0}}$ and $\Delta(\sigma):=\sigma\otimes\sigma$, we have $\Delta(\mathcal{K})=\mathcal{K}\otimes\mathcal{K}$. This is the $q=1$ analog of [BalagovicKolb-qKmatrix, Property (K3’)].
@@ C.4. Construction of Dynamical $K$-matrix
ˇ$\check{\mathcal{R}}_{V,W}(\lambda):=J_{W,V}(\lambda)^{-1}\circ P\circ J_{V,W}(\lambda):V\otimes W\rightarrow W\otimes V$
@@
By naturality of the boundary fusion operator, we have
$(\mathrm{id}_{X}\otimes P)G_{X,W\otimes V}=G_{X,W\otimes V}(\mathrm{id}_{X}\otimes P),$
therefore, the previous equality is equivalent to
@@
which by Lemma C.11 is equivalent to Proposition C.12. ∎
[…]

### v3

@@ Appendix A Technical lemmas
Denote by $D_{i}$ the operator $D_{i}f](\mu):=\xi_{i}(\mu)f_{i}(\mu+2)$. For any rational functions $f_{1}(\mu),\ldots,f_{n}(\mu)$, consider the matrix
@@
We prove the statement by induction. The case $n=1$ is obvious. For the induction step, assume that $f_{1}(\lambda)\neq 0$; if there is $j$ such that $\xi_{j}(\mu)=\xi_{1}(\mu)$ then assume $j=2$. Let us divide the $i$-th row of $A$ by $D_{1}^{i-1}f_{1}](\lambda)$ for every $i$. The matrix becomes
@@
where $\tilde{D}_{i}f](\lambda):=\frac{\xi_{i}(\lambda)}{\xi_{1}(\lambda)}f(\lambda+2)$ and $\tilde{f}_{i}(\lambda):=\frac{f_{i}(\lambda)}{f_{1}(\lambda)}$, and its determinant is still zero. Additionally, let us subtract (i1)(i-1)-th row from the $i$-th one for every $i$. Then we get the matrix
@@
for every $k$ and sufficiently positive $\mu$. In particular, the system given by the matrix (A.1) with the same meaning of notations has a nonzero solution, hence its determinant zero for infinitely many $\mu$. Since the determinant is a rational function in $\mu$, it is zero for all $\mu$. Therefore, by ˜A.1, we get that either one of the functions $f_{i}(\mu)$ is zero or there are $i,j$ such that $\xi_{i}(\mu)=\xi_{j}(\mu)$ (equivalently, $x_{i}=-x_{j}$) and $f_{i}(\mu)=c_{i}f_{j}(\mu)$. In the former case we can apply the induction assumption. As for the latter, denote by
@@
for all sufficiently positive integral $\mu$; now there are n1n-1 rational functions. Using property (7.8), we get for positive integral $\mu$
@@ Appendix B Calculations in rank two
In this appendix, we show the braid relation Theorem˜7.18.
Let $\mathfrak{g}$ be a simple Lie algebra of rank $2$, i.e., $\mathfrak{sl}_{3}$, $\mathfrak{sp}_{4}$, or $\mathfrak{g}_{2}$. Let $\alpha$ (resp. $\beta$) be the short root (resp. the long root), and $\omega_{1}$ (resp. $\omega_{2}$) be the dual fundamental weight. Introduce coordinates on $\mathfrak{h}^{*}$ by $\lambda=\lambda_{1}\omega_{1}+\lambda_{2}\omega_{2}$.
@@
is satisfied if and only if the operators $\{A_{\alpha}^{\mathfrak{k}}(\lambda)\}$ satisfy the shifted one of Theorem˜7.18.
@@
To prove Proposition˜3.13 and Theorem˜7.18, we perform explicit calculations to verify that the operators $R_{\alpha}$ and $R_{\beta}$ satisfy the braid relations in specific integrable $\mathfrak{k}$-modules $\{X_{i}\}$, such that any other irreducible integrable module is a direct $\mathfrak{k}$-summand of $X_{i}\otimes V$ for suitable $i$ and a finite-dimensional $\mathfrak{g}$-representation $V$. These calculations, along with arguments in the proof of Proposition˜3.13 and Theorem˜7.18, imply the braid relations in general integrable $\mathfrak{k}$-modules.
We will also prove ˜7.24 case-by-case. Denote by
@@ B.1.1. Braid relation in special cases
Recall the tensor property $\Delta\tilde{s}^{\mathfrak{k}}_{\alpha}=\tilde{s}^{\mathfrak{k}}_{\alpha}\otimes\tilde{s}^{\mathfrak{g}}_{\alpha}$. By the same arguments as in the proof of Proposition˜3.13 or Theorem˜7.18, it is enough to show this relation for $X=V(1)$ and $V$ an arbitrary $\mathfrak{sl}_{3}$-representation $V$. For the latter, it is enough to show it for the vector representation $\mathbb{C}^{3}$, as any $\mathfrak{sl}_{3}$-representation is a subrepresentation of a suitable tensor product of $\mathbb{C}^{3}$. Both cases can be verified directly.
@@ B.2. Type $C_{2}$
We can choose a basis of $\mathfrak{g}$ satisfying the assumptions of Section˜4 as follows:
@@ B.2.1. Braid relation in special cases
The proof of Proposition˜3.13 is similar to the dynamical case.
B.2.2. Proof of ˜7.24
@@
Therefore, the dynamical braid relation (B.9) for $S$-operators follows from the non-dynamical one of Proposition˜3.13.
@@ B.3. Type $G_{2}$
One can check that $\{J_{i}\}$ and $\{J^{\prime}_{i}\}$ satisfy the $\mathfrak{so}_{3}$-relations (B.5) and $J_{i},J^{\prime}_{k}]=0$ for every $i,k$. In particular, we obtain that $\mathfrak{k}:=\mathfrak{g}_{2}^{\theta}=\mathfrak{so}_{3}\oplus\mathfrak{so}_{3}$. In these generators, we have
@@ B.3.1. Braid relation in special cases
Then we proceed as in Section˜B.1 for both dynamical braid relation of Theorem˜7.18 and non-dynamical one Proposition˜3.13.
B.3.2. Proof of ˜7.24
@@
Denote by $b_{\alpha}^{\mathfrak{so}_{3}},b_{\beta}^{\mathfrak{so}_{3}}$ the type $A_{2}$ generators of Section˜B.1 and by $\Delta$ the coproduct on $\mathfrak{so}_{3}$. Then we have
@@ C.1. Motivation from dynamical $R$-matrix
$\check{\mathcal{R}}_{V,W}(\lambda):=J_{WV}(\lambda)^{-1}\circ P\circ J_{VW}(\lambda),$
@@
$\left(\check{\mathcal{R}}_{V,W}(\lambda)\otimes\mathrm{id}_{U}\right)\circ\left(\mathrm{id}_{V}\otimes\check{\mathcal{R}}_{U,W}(\lambda-h|_{V})\right)\circ\left(\check{\mathcal{R}}_{U,V}(\lambda)\otimes\mathrm{id}_{W}\right).$
@@ C.2. Background on quantum $K$-matrix
If we write $\check{\mathcal{R}}_{q}:=P\circ\mathcal{R}_{q}$, then the $2$-tensor reflection equation becomes
@@
Recall that $w_{0}$ permutes the simple roots, inducing an automorphism of the Dynkin diagram. Let $\sigma$ be the order two automorphism [^10] In general, $\sigma=\theta\theta_{0}$, where $\theta$ is the diagram automorphism in the data of the Satake diagram and $\theta_{0}$ is the automorphism of the Dynkin diagram of $\mathfrak{g}$ induced by $w_{0}$, in this paper we concentrate on the split case – so $\theta=\mathrm{id}$ and $\sigma=\theta_{0}$. of $U_{q}(\mathfrak{g})$ such that $\sigma(E_{\alpha})=E_{-w_{0}(\alpha)}$ and $\sigma(F_{\alpha})=F_{-w_{0}(\alpha)}$ for $\alpha\in R$. The $1$-tensor universal $K$-matrix satisfies the following [Kolb-Kmatrixbraided, equation 1.1]
@@ C.3. Definition of $q=1$ $K$-matrix
Let $w_{0}$ denote the longest element in $W$. The operator $w_{0}$ permutes simple roots, inducing an automorphism of the Dynkin diagram [^11] This automorphism may be $\mathrm{id}$, which is the case if $1\in W\subset GL(\mathfrak{h}^{*})$. which we denote by $\sigma$. The Serre-presentation of $\mathfrak{g}$ implies that $\sigma$ determines an associative algebra automorphism of $\mathrm{U}(\mathfrak{g})$, denoted $\mathrm{Ad}_{\sigma}$ such that $\mathrm{Ad}_{\sigma}(x_{\alpha_{i}})=x_{\alpha_{\sigma(i)}}$, for $x\in\{f,h,e\}$ and $\alpha_{i}\in R$.
@@
As in Remark˜3.2, let $G$ be the simply connected group with the Lie algebra $\mathfrak{g}$. For $g\in G$, we write $\mathrm{Ad}_{g}$ for the automorphism of $\mathfrak{g}$ defined by $x\mapsto gxg^{-1}$. The operators $\widetilde{w}$ can be viewed as elements in $G$ and thus act by conjugation on $\mathfrak{g}$. In particular, one has
@@
When $q=1$, we have $\mathcal{R}_{q=1}=1\otimes 1$ and $\check{\mathcal{R}}_{q=1}=P$. Therefore both sides of the $2$-tensor reflection equation become $\mathcal{K}_{q=1}\otimes\mathcal{K}_{q=1}$. Since $\Delta(\widetilde{w_{0}})=\widetilde{w_{0}}\otimes\widetilde{w_{0}}$ and $\Delta(\sigma):=\sigma\otimes\sigma$, we have $\Delta(\mathcal{K})=\mathcal{K}\otimes\mathcal{K}$. This is the $q=1$ analog of [BalagovicKolb-qKmatrix, Property (K3’)].
@@ C.4. Construction of Dynamical $K$-matrix
$\check{\mathcal{R}}_{V,W}(\lambda):=J_{W,V}(\lambda)^{-1}\circ P\circ J_{V,W}(\lambda):V\otimes W\rightarrow W\otimes V$
@@
Using naturality of $G$, one can show the previous equality is equivalent to
@@
which by Lemma C.11 and naturality of $G$ is equivalent to Proposition C.12. ∎